Splish Co. is a full-service manufacturer of surveillance equipment. Customers can purchase any combination of equipment, installation services, and training as part of Splish’s security services. Thus, each of these performance obligations are separate with individual standalone selling prices. Laplante Inc. purchased cameras, installation, and training at a total price of $82,400. Estimated standalone selling prices of the equipment, installation, and training are $82,800, $6,440, and $2,760, respectively. How should the transaction price be allocated to the equipment, installation, and training?
Answer #1
Stand alone price Percentage of Stand alone price Allocated cost
Equipment 82,800 82,800/92,000 = 90% 82,400 x 90% = $74,160
Installation 6,440 6,440/92,000 = 7% 82,400 x 7% = $5,768
Training 2,760 2,760/92,000 = 3% 82,400 x 3% = $2,472
Total $92,000 100% $82,400
Allocated cost
Equipment $74,160
Installation $5,768
Training $2,472

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.